OpenStudy (anonymous):

Consider the line y=2/3x -6 Find the equation of the line that is perpendicular to this line and passes through the point (4,6) Find the equation of the line that is parallel to this line and passes through the point (4,6)

OpenStudy (anonymous):

y = 2/3 x - 6 has a slope of 2/3 A line that is perpendicular has a slope which is the negative inverse of the slope of the original line. In this case, the negative inverse is -3/2. The point given is (4,6). y=mx+b 6=(-3/2)(4)+b 6=-12/2 + b 6=-6+b b=12 Line perpendicular is: y = (-3/2)x + 12 A line that is parallel shares the same slope as the original line, 2/3. The point given is (4,6). y=mx+b 6=(2/3)(4)+b 6=8/3+b 18/3 = 8/3+b b = 10/3 Line parallel is: y = (2/3)x + 10/3

OpenStudy (anonymous):

both incorrect

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

No. Those answers are perfect. You must have entered them incorrectly.

OpenStudy (mertsj):

Read the directions as to how they are to be entered and follow them.